[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Jun 03, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sun Jun 4 04:52:45 UTC 2017


HSU Shift Log for Sun Jun 04 2017 00:52:38 GMT-0400 (EDT)

Shift: R

Forecaster: R Pasch

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Our COMCAST cable TV was cut off. Roberto Garcia  
(WFO) called them and they said it was because we didn't pay out bill!  
Doubtful this matter will be resolved over the weekend.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170604/a37e524f/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list